OpenStudy (anonymous):

What is the value of x? A. 4 B. 5 C. 8 D. 10

OpenStudy (anonymous):

it would depend on the equation..

OpenStudy (anonymous):

It's an equilateral triangle, so the sides have to be equal. So just fill in x with one of the numbers until they are equal.

OpenStudy (anonymous):

Set the two sides equal to each other and solve for x

OpenStudy (anonymous):

\[5(5)-9=?\]\[3(5)+1=?\]

OpenStudy (anonymous):

16

OpenStudy (anonymous):

5x-9 = 3x +1 2x=10 x=5

OpenStudy (anonymous):

thanks for the help